OpenStudy (anonymous):

graph y= -2x^2+6 describe what you see

OpenStudy (mathmale):

compare your y= -2x^2+6 to y= ax^2 + bx + c (which is one of the standard forms of the equation of a common conic section) Can you recognize this second equation? Note that in your function, a=-2, b+0 and c=6. What would the graph look like?